# BS-LDM: Effective Bone Suppression in High-Resolution Chest X-Ray Images with Conditional Latent Diffusion Models

### v3

@@ 3.1 Data Preparation and Preprocessing
We compiled a dataset of 831 paired posterior-anterior CXR and DES soft tissue images. These images were acquired using a digital radiography system with a dual-exposure DES unit (Discovery XR656, GE Healthcare) from our partner hospital. Originally archived in 14-bit DICOM format, the images were converted to PNG format to simplify processing procedures. Each image measures $2021\times 2021$ pixels, with pixel sizes ranging from 0 to 0.1943 mm.
From this collection, we excluded 13 image pairs due to operational errors, severe motion artifacts, and conditions like pleural effusions and pneumothorax, which could interfere with analysis. The refined dataset, named SZCH-X-Rays, consists of 818 image pairs. These were divided into training, validation, and test sets in an 8:1:1 ratio. We enhanced local contrast in the images using contrast-limited adaptive histogram equalization.
@@ 3.3 Evaluation Metrics
Four metrics are used to assess the quality of bone suppression: the Bone Suppression Ratio (BSR) [33], Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR) [34], and Learned Perceptual Image Patch Similarity (LPIPS) [35].
The BSR evaluates how effectively bone suppression is applied in CXR images. It measures the difference between the estimated soft tissue image, denoted as $\hat{\mathbf{S}}$, and the actual soft tissue image, $\mathbf{S}$, while considering the corresponding bone image $\mathbf{B}$. The BSR formula is expressed as follows:
$BSR=1-\sum_{i=1}^{N}(\mathbf{S}_{i}-\hat{\mathbf{S}}_{i})^{2}/\sum_{i=1}^{N}\mathbf{B}_{i}^{2},$ (13)
where a BSR value of 1 indicates ideal performance.
In addition to BSR, we employ two common pixel-level metrics, MSE and PSNR, which assess the pixel-level discrepancies between the generated and actual soft tissue images.
To further evaluate the generation of detail from a perceptual standpoint, we utilize LPIPS. This metric measures perceptual similarity between two images by comparing features extracted from a pre-trained deep neural network. LPIPS is calculated as follows:
$LPIPS=\frac{1}{N\cdot M\cdot F}\sum_{i=1}^{N}\sum_{j=1}^{M}\sum_{f=1}^{F}\left(\phi_{i,j,f}(\mathbf{S})-\phi_{i,j,f}(\hat{\mathbf{S}})\right)^{2},$ (14)
@@
3.4 Comparisons
To ensure an unbiased evaluation, we compared our proposed model with the leading ones recently highlighted in the literature. Specifically, we assessed our model against: an autoencoder-based model [12], the GAN-based MCA-Net [14], the ResNet-based ResNet-BS [16], and the diffusion model-based BS-Diff [21].
@@ 3.4.1 Evaluation on the SZCH-X-Rays Dataset
[…]
@@ 3.4.2 Evaluation on the JSRT Dataset
[…]
Figure 6: Visualization of ablation studies of offset noise and the temporal adaptive thresholding strategy on BS-LDM, with histograms given to visualize the pixel intensity distribution more intuitively.
Table 1: Comparison on SZCH-X-Rays and JSRT datasets with the state-of-the-art methods using BSR, MSE, PSNR and LPIPS as evaluation metrics.
@@
SZCH-X-Rays Gusarev et al. [12] 0.944 $\pm$ 0.018 0.0008 $\pm$ 0.0002 31.900 $\pm$ 3.449 0.220 $\pm$ 0.012
MCA-Net [14] 0.945 $\pm$ 0.019 0.0007 $\pm$ 0.0002 32.177 $\pm$ 2.912 0.118 $\pm$ 0.011
ResNet-BS [16] 0.956 $\pm$ 0.021 0.0009 $\pm$ 0.0003 31.950 $\pm$ 3.097 0.078 $\pm$ 0.008
BS-Diff [21] 0.971 $\pm$ 0.022 0.0008 $\pm$ 0.0004 31.057 $\pm$ 3.585 0.191 $\pm$ 0.015
BS-LDM (Ours) 0.976 $\pm$ 0.018 0.0006 $\pm$ 0.0003 33.224 $\pm$ 3.577 0.051 $\pm$ 0.016
JSRT Gusarev et al. [12] 0.865 $\pm$ 0.041 0.0009 $\pm$ 0.0002 31.216 $\pm$ 2.255 0.193 $\pm$ 0.017
MCA-Net [14] 0.892 $\pm$ 0.039 0.0009 $\pm$ 0.0002 33.735 $\pm$ 1.578 0.148 $\pm$ 0.017
ResNet-BS [16] 0.903 $\pm$ 0.038 0.0009 $\pm$ 0.0003 32.009 $\pm$ 1.653 0.139 $\pm$ 0.020
BS-Diff [21] 0.902 $\pm$ 0.032 0.0008 $\pm$ 0.0003 33.937 $\pm$ 2.465 0.147 $\pm$ 0.024
BS-LDM (Ours) 0.908 $\pm$ 0.029 0.0007 $\pm$ 0.0002 34.312 $\pm$ 2.176 0.049 $\pm$ 0.018
@@
3.5 Ablation Study
3.5.1 Offset Noise Effect
In examining the effectiveness of offset noise, we found that its application resulted in generated soft tissue images with mean values more aligned with real images. This adjustment led to a reduction in MSE by at least 92.5%, an improvement in PSNR by at least 0.897 dB, and a decrease in LPIPS by at least 3.8% compared to images generated without offset noise, across both datasets. Fig. 6 highlights that without offset noise, the brightness of the generated images was significantly misaligned.
3.5.2 Temporal Adaptive Thresholding Effect
@@
* $\lambda_{L1}$, $\lambda_{Qua}$, $\lambda_{Per}$ and $\lambda_{Adv}$, from left to right.
@@
* $\lambda_{L1}$, $\lambda_{Qua}$, $\lambda_{Per}$ and $\lambda_{Adv}$, from left to right.
@@
Junior Radiologist Precision ($\uparrow$) Recall ($\uparrow$) F1 Score ($\uparrow$)
@@
Senior Radiologist Precision ($\uparrow$) Recall ($\uparrow$) F1 Score ($\uparrow$)
@@ 3.6.1 Hyperparameter Analysis on the ML-VQGAN Loss
We conducted a detailed analysis for each dataset to further examine the effects of the multi-level hybrid loss function of ML-VQGAN under varying hyperparameter settings. The results, as depicted in Table 3 and Table 4, reveal that the BS-LDM framework achieves optimal performance when the weight parameters $\lambda_{L1}$, $\lambda_{Qua}$, $\lambda_{Per}$, and $\lambda_{Adv}$ are set to 1, 1, 0.001, and 0.01, respectively.
@@ 3.7.2 Diagnostic Utility Assessment
The diagnostic value of soft tissue imaging was independently evaluated by two radiologists with 6 and 11 years of experience, following the X-ray diagnosis standard [38]. This analysis employed the SZCH-X-Rays dataset for bone suppression, using computed tomography to confirm lesions, which included common lung diseases such as inflammation, tuberculosis, and masses or nodules. Out of 818 data pairs assessed, 79 pairs contained one or more of these lesions. The radiologists independently evaluated both conventional CXR and the soft tissue images generated by our model. Table 6 presents the results, highlighting key metrics such as Precision, Recall, and the F1 Score. The findings suggest that the soft tissue images produced by BS-LDM enable more thorough and accurate lesion diagnosis compared to CXR images, thereby confirming its high clinical diagnostic value.
@@ 3.8 Limitations
While BS-LDM demonstrates exceptional performance in suppressing bone structures and generating high-resolution CXR images, several limitations remain, and there are opportunities for further enhancements. For example, incorporating more advanced denoising samplers [32, 39] could help reduce the number of sampling steps needed. Additionally, adopting a mask-based approach might improve detail control across different regions of the CXR images. Another potential improvement involves replacing the noise estimator network with alternatives, such as Transformer models [40] or Structured Space Models (SSM) [41], which could enhance the model’s flexibility and scalability. Finally, the applicability of our method in downstream computerized analyses [42, 43] remains to be confirmed. In our future work, we plan to address these limitations and explore potential improvements.

### v5

@@ 3.1 Data Preparation and Preprocessing
We compiled a dataset of 831 paired posterior-anterior CXR and DES soft tissue images. These images were acquired using a digital radiography system with a dual-exposure DES unit (Discovery XR656, GE Healthcare) from our partner hospital. Originally archived in 14-bit DICOM format, the images were converted to PNG format to simplify processing procedures. Each image measures $2021\times 2021$ pixels, with pixel sizes ranging from 0 to 0.1943 mm. From this collection, we excluded 13 image pairs due to operational errors, severe motion artifacts, and conditions like pleural effusions and pneumothorax, which could interfere with analysis. The refined dataset, named SZCH-X-Rays, consists of 818 image pairs. These were divided into training, validation, and test sets in an 8:1:1 ratio. We enhanced local contrast in the images using contrast-limited adaptive histogram equalization.
@@ 3.2 Implementation Details
Figure 5: Bone suppression and pseudo-color zoomed-in views of BS-LDM compared with state-of-the-art approaches on SZCH-X-Rays (top) and JSRT (bottom) datasets.
@@ 3.3 Evaluation Metrics
We evaluate the quality of bone suppression using four metrics: Bone Suppression Ratio (BSR) [35], Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR) [36], and Learned Perceptual Image Patch Similarity (LPIPS) [37]. These metrics are applied in both comparative analyses and ablation studies.
The BSR evaluates how effectively bone suppression is applied in CXR images. It measures the difference between the estimated soft tissue image, denoted as $\hat{S}$, and the actual soft tissue image, $S$, while considering the corresponding bone image $B$. The BSR formula is expressed as follows:
$BSR=1-\sum_{i=1}^{N}(S_{i}-\hat{S}_{i})^{2}/\sum_{i=1}^{N}B_{i}^{2},$ (12)
where a BSR value of 1 indicates ideal performance. In addition to BSR, we employ two common pixel-level metrics, MSE and PSNR, which assess the pixel-level discrepancies between the generated and actual soft tissue images. To further evaluate the generation of detail from a perceptual standpoint, we utilize LPIPS. This metric measures perceptual similarity between two images by comparing features extracted from a pre-trained deep neural network. LPIPS is calculated as follows:
$LPIPS=\frac{1}{N\cdot M\cdot F}\sum_{i=1}^{N}\sum_{j=1}^{M}\sum_{f=1}^{F}\left(\phi_{i,j,f}(S)-\phi_{i,j,f}(\hat{S})\right)^{2},$ (13)
@@
To evaluate the downstream diagnostic value of our bone suppression method, we employ five metrics: Precision, Sensitivity, F1 Score, Accuracy, and Specificity. Precision quantifies the proportion of true positives among all positive classifications, minimizing false positives. Sensitivity (Recall) measures the method’s ability to detect true pathological findings in bone-suppressed images, which is particularly important for disease screening applications. The F1 Score balances Precision and Sensitivity, providing a comprehensive metric that mitigates potential biases arising from class imbalances. Accuracy reflects the overall correctness of diagnostic predictions across all categories, ensuring clinical reliability. Lastly, Specificity evaluates the method’s ability to correctly identify normal anatomical structures, minimizing false alarms and preserving accuracy in healthy case interpretations.
3.4 Comparison Results
To ensure a fair and comprehensive evaluation, we compared our proposed model against several state-of-the-art approaches recently highlighted in the literature. Specifically, our assessment included comparisons with an autoencoder-based model [12], the GAN-based MDA-GAN [19], method by Rani et al. [18] and MCA-Net [14], the ResNet-based ResNet-BS [16], the U-Net-based xU-NetFullSharp [20], and the diffusion model-based BS-Diff [24].
@@ 3.4 Comparison Results
Figure 6: Inference efficiency comparison of our method and other diffusion-based methods on the SZCH-X-Rays dataset on a single Nvidia A100 80G GPU.
@@ 3.4.1 Evaluation on the SZCH-X-Rays Dataset
In our SZCH-X-Rays dataset, the proposed method demonstrates superior performance across all evaluation metrics compared to existing approaches. Specifically, it achieves improvements in BSR by 1.6% ($\uparrow$), MSE by 16.7% ($\downarrow$), PSNR by 1.037 dB ($\uparrow$), and LPIPS by 34.6% ($\downarrow$). The reduction in LPIPS, exceeding 34.6%, underscores the enhanced retention of critical lung details, as illustrated in Fig. 5. Among the evaluated methods, the autoencoder-based model and BS-Diff perform poorly across most metrics, largely due to significant texture loss. While MDA-GAN and xU-NetFullSharp exhibit relatively better detail retention, they fall short of the proposed BS-LDM method due to minor inconsistencies in brightness and detail preservation. Notably, all methods achieve effective bone suppression, with a minimum BSR of 94.4%. However, BS-LDM surpasses them with a BSR of 97.6%. Additionally, as shown in Fig. 6, BS-LDM demonstrates efficient processing with an inference time approximately 77.7% of that required by BS-Diff and DDPM, further solidifying its practicality for clinical applications.
@@ 3.4.2 Evaluation on the JSRT Dataset
On the public JSRT dataset, our method consistently outperforms all other approaches across key metrics, achieving improvements in BSR by 2.1% ($\uparrow$), MSE by 11.3% ($\downarrow$), PSNR by 2.315 dB ($\uparrow$), and LPIPS by 31.0% ($\downarrow$). As illustrated in Fig. 5, BS-LDM produces sharper and more detailed outputs compared to BS-Diff and MCA-Net. However, the distinctions are less pronounced than those observed in the SZCH-X-Rays dataset, likely due to the inherent blurring present in the JSRT dataset. Quantitatively, BS-LDM delivers significantly lower LPIPS scores on the JSRT dataset, reflecting enhanced perceptual quality and better preservation of soft tissue details. By contrast, the autoencoder-based model, MCA-Net, Rani et al., and BS-Diff demonstrate mediocre performance across most metrics, which can be attributed to minor pixel-level luminance inconsistencies and a lack of fine detail retention.
Figure 7: Visualization of ablation studies on BS-LDM. The upper row shows generated soft tissue images under different conditions, while the lower row presents corresponding pixel intensity histograms, where rightward shift indicates brighter images, and wider spread indicates higher contrast.
Table 1: Comparison on SZCH-X-Rays and JSRT datasets with the state-of-the-art methods using four evaluation metrics.
@@
SZCH-X-Rays Gusarev et al. [12] 0.944 $\pm$ 0.018 0.00084 $\pm$ 0.00023 31.900 $\pm$ 3.449 0.220 $\pm$ 0.012
MCA-Net [14] 0.945 $\pm$ 0.019 0.00073 $\pm$ 0.00021 32.177 $\pm$ 2.912 0.118 $\pm$ 0.011
ResNet-BS [16] 0.956 $\pm$ 0.021 0.00085 $\pm$ 0.00034 31.850 $\pm$ 3.097 0.078 $\pm$ 0.008
Rani et al. [18] 0.955 $\pm$ 0.031 0.00082 $\pm$ 0.00039 31.961 $\pm$ 2.924 0.084 $\pm$ 0.022
MDA-GAN [19] 0.957 $\pm$ 0.025 0.00075 $\pm$ 0.00026 32.180 $\pm$ 3.077 0.128 $\pm$ 0.026
xU-NetFullSharp [20] 0.950 $\pm$ 0.014 0.00072 $\pm$ 0.00038 32.187 $\pm$ 3.747 0.125 $\pm$ 0.024
BS-Diff [24] 0.961 $\pm$ 0.022 0.00077 $\pm$ 0.00043 32.181 $\pm$ 3.296 0.191 $\pm$ 0.015
BS-LDM (Ours) 0.976 $\pm$ 0.018 0.00060 $\pm$ 0.00032 33.224 $\pm$ 3.577 0.051 $\pm$ 0.016
JSRT Gusarev et al. [12] 0.865 $\pm$ 0.041 0.00094 $\pm$ 0.00024 30.006 $\pm$ 2.234 0.193 $\pm$ 0.017
MCA-Net [14] 0.892 $\pm$ 0.039 0.00090 $\pm$ 0.00023 31.215 $\pm$ 1.576 0.148 $\pm$ 0.017
ResNet-BS [16] 0.903 $\pm$ 0.038 0.00089 $\pm$ 0.00029 31.239 $\pm$ 1.639 0.139 $\pm$ 0.020
Rani et al. [18] 0.896 $\pm$ 0.037 0.00087 $\pm$ 0.00031 31.435 $\pm$ 2.246 0.081 $\pm$ 0.016
MDA-GAN [19] 0.894 $\pm$ 0.035 0.00085 $\pm$ 0.00034 31.671 $\pm$ 2.277 0.082 $\pm$ 0.018
xU-NetFullSharp [20] 0.900 $\pm$ 0.055 0.00084 $\pm$ 0.00045 31.868 $\pm$ 3.879 0.071 $\pm$ 0.017
BS-Diff [24] 0.902 $\pm$ 0.032 0.00080 $\pm$ 0.00033 31.997 $\pm$ 2.555 0.147 $\pm$ 0.024
BS-LDM (Ours) 0.922 $\pm$ 0.028 0.00071 $\pm$ 0.00024 34.312 $\pm$ 2.176 0.049 $\pm$ 0.018
@@
[1, 0, $10^{-3}$, $10^{-2}$] 0.00390 27.130 0.185
[0, 1, $10^{-3}$, $10^{-2}$] 0.00128 28.743 0.062
$\star$ $\lambda_{L1}$, $\lambda_{Qua}$, $\lambda_{Per}$ and $\lambda_{Adv}$, from left to right.
@@
[1, 0, $10^{-3}$, $10^{-2}$] 0.00237 27.636 0.172
[0, 1, $10^{-3}$, $10^{-2}$] 0.00151 28.392 0.059
$\star$ $\lambda_{L1}$, $\lambda_{Qua}$, $\lambda_{Per}$ and $\lambda_{Adv}$, from left to right.
@@
3.5 Ablation Study
3.5.1 The Effect of Offset Noise
[…]
3.5.2 The Effect of Temporal Adaptive Thresholding
@@
Junior Radiologist Precision ($\uparrow$) Sensitivity ($\uparrow$) F1 Score ($\uparrow$)
@@
Senior Radiologist Precision ($\uparrow$) Sensitivity ($\uparrow$) F1 Score ($\uparrow$)
@@
Table 7: Automated downstream evaluation of BS-LDM on Shenzhen chest X-ray using three classification models.
Modal Method Precision ($\uparrow$) Sensitivity ($\uparrow$) F1 Score ($\uparrow$) Accuracy ($\uparrow$) Specificity ($\uparrow$)
CXR AlexNet [38] 0.816 $\pm$ 0.004 0.838 $\pm$ 0.068 0.826 $\pm$ 0.028 0.821 $\pm$ 0.026 0.803 $\pm$ 0.017
DenseNet [39] 0.824 $\pm$ 0.014 0.926 $\pm$ 0.051 0.872 $\pm$ 0.032 0.862 $\pm$ 0.026 0.795 $\pm$ 0.035
ResNet [40] 0.816 $\pm$ 0.077 0.912$\pm$ 0.051 0.861 $\pm$ 0.065 0.851 $\pm$ 0.069 0.788 $\pm$ 0.087
Soft Tissue AlexNet [38] 0.855 $\pm$ 0.015 0.951 $\pm$ 0.017 0.898 $\pm$ 0.016 0.891 $\pm$ 0.017 0.828 $\pm$ 0.017
DenseNet [39] 0.833 $\pm$ 0.020 0.956 $\pm$ 0.051 0.890 $\pm$ 0.033 0.881 $\pm$ 0.034 0.803 $\pm$ 0.017
ResNet [40] 0.841 $\pm$ 0.035 0.941 $\pm$ 0.068 0.888 $\pm$ 0.050 0.881 $\pm$ 0.052 0.818 $\pm$ 0.035
@@ 3.6.1 Hyperparameter Analysis on the ML-VQGAN Loss
We conducted a detailed analysis for each dataset to further examine the effects of the multi-level hybrid loss function of ML-VQGAN under varying hyperparameter settings. The results, as depicted in Table 3 and Table 4, reveal that the BS-LDM framework achieves optimal performance when the weight parameters $\lambda_{L1}$, $\lambda_{Qua}$, $\lambda_{Per}$, and $\lambda_{Adv}$ are set to 1, 1, 0.001, and 0.01, respectively. The removal of the $\mathcal{L}_{Qua}$ results in a noticeable decline across all evaluation metrics. Even if the primary reconstruction loss is replaced with $\mathcal{L}_{Per}$, the absence of the $\mathcal{L}_{L1}$ leads to a significant reduction in pixel-level performance, though the impact on LPIPS is relatively minor. Moreover, when both the $\mathcal{L}_{Per}$ and $\mathcal{L}_{Adv}$ are omitted, the generated images exhibit a lack of texture detail. This deficiency is reflected in a consistent drop across all metrics, underscoring the importance of these components for achieving high-quality results.
@@ 3.7.2 Diagnostic Utility Assessment
The diagnostic value of soft tissue imaging was independently evaluated by two radiologists with 6 and 11 years of experience, following the X-ray diagnosis standard [43]. This analysis employed the SZCH-X-Rays dataset for bone suppression, using computed tomography to confirm lesions, which included common lung diseases such as inflammation, tuberculosis, and masses or nodules. Out of 818 data pairs assessed, 79 pairs contained one or more of these lesions. The radiologists independently evaluated both conventional CXR and the soft tissue images generated by our model. Table 6 summarizes the evaluation results, focusing on key diagnostic metrics such as Precision, Sensitivity, and F1 Score. The findings reveal that the soft tissue images generated by BS-LDM significantly enhance diagnostic performance for both junior and senior radiologists compared to conventional CXR images. Notably, the diagnostic accuracy achieved by junior radiologists using the generated soft tissue images closely matches the performance of senior radiologists working with CXR images. These results underscore the potential of BS-LDM to improve lesion detection and diagnosis, offering a more comprehensive and accurate assessment. This highlights its substantial value in clinical applications, where it could serve as an effective tool to augment diagnostic capabilities across varying levels of expertise.
3.8 Automated Downstream Evaluation
To objectively evaluate the effectiveness of our method for lung disease diagnosis, we performed an automated downstream analysis on the widely used public Shenzhen chest X-ray dataset [44]. This dataset includes 336 radiographs of tuberculosis cases and 326 normal cases. Using the BS-LDM model trained on SZCH-X-Rays, we generated corresponding soft tissue images for the Shenzhen chest X-ray dataset. Both the original CXR images and the generated soft tissue images were input into lung disease prediction systems based on AlexNet [38], DenseNet [39], and ResNet [40] for evaluation. The quantitative results, presented in Table 7, demonstrate consistent performance improvements across key metrics, including Precision, Sensitivity, F1 Score, Accuracy, and Specificity. Notably, when using soft tissue images generated by BS-LDM, Sensitivity improved by 13.5%, 3.24%, and 3.18% in AlexNet, DenseNet, and ResNet models, respectively. These enhancements highlight the capability of our method to augment downstream diagnostic performance in CXR imaging, further supporting its clinical utility in improving lung disease detection.
@@ 3.9 Limitations
While BS-LDM has shown remarkable capability in producing high-resolution soft tissue images from CXR data, its application in clinical practice remains limited. To enable successful integration into existing Picture Archiving and Communication Systems (PACS), the following key challenges have to be addressed:
  • Although BS-LDM has achieved notable improvements in inference efficiency, surpassing other diffusion-based methods by 28.6%, as shown in Fig. 6, there remains room for further optimization. To enhance its performance, we intend to investigate the incorporation of advanced denoising samplers, such as DPM [34] and LCM [45]. These approaches have the potential to reduce the number of sampling steps, thereby increasing the model’s overall speed and efficiency.
  • Processing high-resolution images often demands substantial memory, which may restrict the practical use of BS-LDM in clinical settings with limited computational resources. To overcome this challenge, we are investigating architectural advancements that aim to reduce memory consumption while preserving performance. Promising approaches involve incorporating Transformer models [46] or Structured Space Models [47] as noise estimator networks. These innovations could enable a more memory-efficient architecture that retains the precise detail required for medical imaging, thereby enhancing the model’s suitability for deployment across clinical environments with diverse resource constraints.
  • Ensuring compatibility with standard medical image formats, such as DICOM, is essential for the seamless integration of BS-LDM into PACS. To address this, we are evaluating BS-LDM’s ability to handle DICOM-formatted data during both pre-processing and post-processing stages, thereby streamlining its incorporation into routine clinical workflows.
  • Developing a user-friendly interface is crucial for integrating BS-LDM seamlessly into existing clinical workflows. To ensure practicality and ease of use, we are collaborating with radiologists to design an intuitive interface that streamlines interaction with the model, promoting greater user adoption and enhancing its clinical utility.
In future research, we aim to address these limitations and advance the clinical application of our method.